OpenStudy (anonymous):

Determine the sum of the infinite geometric series -1-1/sqrt6-1/6+...

OpenStudy (anonymous):

that is a geometric series, S=a+a*q+a*q^2+.... you should first determine the value of q and then you can calculate the summation by a well-known formula

OpenStudy (anonymous):

so would it be -1/(1-1/sqrt6) ?

OpenStudy (anonymous):

see the q=1/sqrt(6), and as the q<1 we may write: S=a/(1-q)=-1/(1-1/sqrt(6)). so you are right

OpenStudy (anonymous):

Okay! so what exactly would the sum be? i'm still a bit confused

OpenStudy (anonymous):

take it easy. you are a sharp student just trust yourself. (your answer is right)

OpenStudy (anonymous):

According to my homework site it is not correct unfortunately

OpenStudy (anonymous):

repeat it once again but I think that is true